OpenStudy (anonymous):

What are the zeros of the function? What are their multiplicities? y=2x(x+1)³

OpenStudy (ranga):

To find zeros, set y = 0 and solve for x.

OpenStudy (anonymous):

how do i solve

OpenStudy (ranga):

2x(x+1)³ = 0 x = 0 or (x+1)³ = 0 x = 0 or x = -1 Therefore, the zeros are: 0 and -1 Since x occurs just once, the multiplicity of 0 is 1. Since (x+1)³ = (x+1)(x+1)(x+1) = 0, the zero at -1 occurs three times. Multiplicity = 3 The zeros are: 0 and -1 and their multiplicities are 1 and 3 respectively.